OpenStudy (anonymous):

Hello guys I need help with solving this combining like terms problem. OK -1/2a+3a-1/2+3/4

OpenStudy (anonymous):

I get 11/4 + a squared

OpenStudy (mathstudent55):

Do you know what like terms are?

OpenStudy (anonymous):

Yes I do

OpenStudy (mathstudent55):

Ok, now look at the expression you have, which terms are like terms?

OpenStudy (anonymous):

Well you have the fractions and the variables

OpenStudy (mathstudent55):

Let me give you an example without variables to explain like terms. Let's say you have 5 oranges and 2 apples. I give you 3 oranges and 4 apples. What do you have now?

OpenStudy (mathstudent55):

Your problem is to combine like terms in this expression: 5 oranges + 2 apples + 3 oranges + 2 apples What is the sum?

OpenStudy (anonymous):

8 oranges + 4 apples but you would have 12 fruits

OpenStudy (mathstudent55):

Yes, you are correct in both statements. In math, though, we keep the variables separate. If you have 5a + 3b + 2a + 4b, you can combine the terms with a together, and you can combine the terms with b together, but you can't combine terms with a and be together. So 5a + 3b + 2a + 4b = 5a + 2a + 3b + 4b = 7a + 7b There is no more you can do.

OpenStudy (mathstudent55):

Now let's get back to your problem.

OpenStudy (mathstudent55):

Like terms have exactly the same variable parts. Terms with a are like terms and can be combined. Terms with no variable are like terms and can be combined. Terms with a and terms without a variable are NOT like terms and CANNOT be combined toghether.

OpenStudy (mathstudent55):

This is your problem, correct? \( \dfrac{1}{2}a+3a-\dfrac{1}{2}+\dfrac{3}{4} \)

OpenStudy (mathstudent55):

The first two terms have the variable a. They are like terms and can be combined together. The last two terms have no variable. They are like terms and can be combined together.

OpenStudy (anonymous):

So is it 7/2 a -1/4?

OpenStudy (anonymous):

But actually the textbook is telling me that 1/2 a is 1/2 * a

OpenStudy (mathstudent55):

Since we have fractions, we need common denominators. \(\dfrac{1}{2}a+3a-\dfrac{1}{2}+\dfrac{3}{4}\) \(= \dfrac{1}{2}a+\dfrac{3}{1}a -\dfrac{1}{2}+\dfrac{3}{4} \) \(= \dfrac{1}{2}a+\dfrac{2}{2} \times \dfrac{3}{1}a -\dfrac{2}{2} \times \dfrac{1}{2}+\dfrac{3}{4} \) \(= \dfrac{1}{2}a+\dfrac{6}{2}a -\dfrac{2}{4}+\dfrac{3}{4} \) \(= \dfrac{7}{2}a +\dfrac{1}{4} \)

OpenStudy (anonymous):

Well I got - 1/4 so yes I was right except for the sign well thank you :-) for helping me mathstudent!

OpenStudy (mathstudent55):

You were close. Good effort. You're welcome.